OpenStudy (anonymous):

help! slope between (-1,2) and (4,4) pleasse help

OpenStudy (anonymous):

Do you mean find the slope of the line that passes through those two points?

OpenStudy (anonymous):

to find slope its \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

ex:(0,3 and 3,0) o-3 3-0 -3 3=1

OpenStudy (anonymous):

HELP WITH WRITING

OpenStudy (anonymous):

y2=4 y1=2 x2=4 x1=-1

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = (4-2)/(4-(-1)) m = (4-2)/(4+1) m = 2/5 So the slope of the line that passes through the two points is 2/5

OpenStudy (anonymous):

If you want to find the slope, you need to use this equation: m = y2 - y1 / x2 - x1 Plugging in the numbers from your question: m = 4 - 2 / 4 - (-1) Simplify. m = 2 / 5 So the slope is 2/5. And blondie16 was correct. ^-^ If you mean something else, I don't know how to help you..